# RDB2G-Bench: A Comprehensive Benchmark for Automatic Graph Modeling of Relational Databases

### v1

@@ 5.2 Benchmark Results
For fair comparisons, we evaluate all baselines under a budget, which limits the number of graph models whose effectiveness (e.g., ground-truth downstream-task performance) can be measured. We track the effectiveness of resulting graph models as the budget increases up to the 5% of the total search space. Full results, omitted here due to space constraints, are provided in Appendix D.
Figure LABEL:fig:benchmark_rank shows the ranks of the baselines, averaged over 12 predictive tasks from RDB2G-Bench and 10 independent runs per baseline (3 runs for the LLM-based method). Based on these results, we derive several key observations:
@@
  • Greedy methods perform comparably to advanced approaches such as Bayesian Optimization and Reinforcement Learning. Especially, Greedy Backward performs best under small budgets, while Greedy Forward outperforms it as the budget increases. Therefore, the choice between them can be guided by the available computational resources or time constraints.
  • In the short run, the LLM-based approach demonstrated strong capabilities in rapidly improving performance through effective multi-turn planning. However, it struggled to devise effective long-term action plans, resulting in limited gains as the budget increased.
@@
Efficiency Gain due to RDB2G-Bench: Our extensive benchmarking is made efficient by RDB2G-Bench, i.e., the precomputed performance metrics for graph models. Without RDB2G-Bench, it is inevitable to evaluate graph models on the fly during the search process, requiring repeated training of predictive models (e.g., graph neural networks). As shown in Figure LABEL:fig:benchmark_time, using RDB2G-Bench reduces evaluation time 606$\times$ from over 800 hours to just 1.38 hours. RDB2G-Bench equally benefits new RDB-to-graph modeling methods, fostering research in this direction.
@@ 6 Conclusions
This study presents RDB2G-Bench, the first benchmark designed to evaluate automatic gr‘aph modeling methods for relational databases (RDBs). For RDB2G-Bench, we precomputed 50k graph models derived from 5 real-world RDBs and 12 predictive tasks, along with their associated downstream performance, runtime, and parameter size.
Analysis of the RDB2G-Bench datasets reveals that more data or additional tables do not necessarily improve predictive performance. Instead, selectively using fewer, relevant tables often leads to better results with improved efficiency. Moreover, the best-performing graph models vary by task, emphasizing the need for intelligent modeling strategies that account for the characteristics of both RDBs and predictive tasks.
Moreover, our extensive benchmark of nine RDB-to-graph modeling strategies reveals substantial room for improvement in this domain, with LLM-based reasoning showing promising potential, despite its current limitations.
RDB2G-Bench, whose datasets and code are publicly available at https://github.com/chlehdwon/RDB2G-Bench, significantly accelerates the evaluation of diverse RDB-to-graph modeling strategies, by up to 606$\times$, thereby facilitating the development of more advanced techniques.
@@ A.1 RelBench Detail
Avito is a leading online advertisement platform, providing a marketplace for users to buy and sell a wide variety of products and services, including real estate, vehicles, jobs, and goods. The Avito Context Ad Clicks dataset on Kaggle is part of a competition aimed at predicting whether an ad will be clicked based on contextual information. This dataset includes user searches, ad attributes, and other related data to help build predictive models.
@@ A.2 Model Implementation
In our implementation, we employ Equation (1) and Equation (2) for the $\mathrm{LocalMP}^{(k)}$ of Row2Node and Row2Edge modeling, respectively, to naturally extend our settings to transformer-based architectures. To fit the GPS model within a single GPU, we reduce the number of sampled neighbors from 128 to 32.
@@ A.3 Hyperparameter Details
Table 2 shows our learning rate choices, which are only tuned for each task. For the other hyperparameter, all tasks used a batch size of 512, 128 dimensions, and the Adam [21] optimizer.
Specifically, neighbor sampling counts for subgraph extraction are set to 128 when all tables are represented as nodes, and 12 when some tables are represented as edges. A lower sampling count is employed when tables are modeled as edges, since the two-layer model effectively captures paths longer than two hops, resulting in a significantly larger graph scale that cannot be processed by a single GPU.
@@ Appendix B Additional Results regarding Our Observations on the RDB2G-Bench Datasets
As shown in Figure 25, while cross-model correlations are relatively low for the two tasks on the rel-f1 dataset (driver-top3 and driver-position) due to high variances in the performances, for the other tasks, cross-model correlations are high. The results for user-repeat, user-visits, and user-ad-visit are presented in Figure 6 in the main paper.
Figure 25: Additional results regarding Obs 5. Spearman correlations between different predictive models—GraphSAGE with sum and mean aggregation (denoted as sum and mean), GIN, and GPS—are generally high (above 0.7) across the remaining tasks.
@@ C.1 Implementation Details of Baselines
Bayesian Optimization. The Bayesian Optimization (BO) baseline employs an iterative search leveraging a surrogate Multi-Layer Perceptron (MLP) model based on the BANANAS [35] implementation. The input embeddings represent graph structures converted into fixed-sized arrays based on selected edges.
@@ D.1 Performance Results on the RDB2G-Bench Datasets
In this section, we provide additional performance details of nine RDB-to-graph modeling methods, supplementing the summary presented in Figure LABEL:fig:benchmark_rank in the main paper. Details are presented in Figure 26.
Figure 26: Performance details of nine RDB-to-graph modeling methods on each of the 12 predictive tasks. The performances are computed under varying budget levels, corresponding to the number of graph models evaluated.
@@ D.2 Analysis of Evaluation Time
In this section, we present a detailed analysis of evaluation time costs for each RDB-to-graph modeling method, which complements the results summarized in Figure LABEL:fig:benchmark_time in the main paper. Details are provided in Table 3.
@@
Random 63803.69 0.59 107072.34
Greedy Forward Search (GF) 20849.53 0.39 53033.08
Greedy Backward Search (GB) 43206.87 0.51 84005.94
Greedy Local Search (GL) 56888.20 0.34 164955.41
Evolutionary Algorithm (EA) 139491.21 132.62 1051.85
Bayesian Optimization (BO) 120166.04 1244.17 96.58
Reinforcement Learning (RL) 150787.19 17.66 8537.55
LLM-based baseline (LLM) 13042.90 3588.22 3.63

### v2

@@ 5.2 Benchmark Results
(a) Rank Averaged over Tasks
(b) Evaluation Cost
[…]
For fair comparisons, we evaluate all baselines under a budget, which limits the number of graph models whose effectiveness (e.g., ground-truth downstream-task performance) can be measured. We track the effectiveness of resulting graph models as the budget increases up to 5% of the total search space. Full results, omitted here due to space constraints, are provided in Appendix D.
Figure 7(a) shows the ranks of the baselines, averaged over 12 predictive tasks from RDB2G-Bench and 10 independent runs per baseline (3 runs for the LLM-based method). In addition, Table 1 shows the details results on the rel-f1 dataset. Based on these results, we derive several key observations:
@@
  • Greedy methods and Random perform comparably to advanced approaches such as Bayesian Optimization and Reinforcement Learning. Especially, Greedy Backward performs best under small budgets, while Bayesian Optimization outperforms it as the budget increases. Therefore, the choice between them can be guided by the available computational resources or time constraints.
  • Value-based optimization algorithms, such as Bayesian Optimization and Greedy Forward, typically improve in effectiveness as performance feedback accumulates. However, under limited compute budgets, more complex algorithms, such as Reinforcement Learning, or algorithms affected by poor initialization, such as Greedy Local, often suffer from unstable exploration, resulting in poor performance.
  • In the short run, the LLM-based approach demonstrated strong capabilities in rapidly improving performance through contextual reasoning and effective multi-turn planning. However, it struggled to devise effective long-term action plans, resulting in limited gains as the budget increased. CoT reasoning shows slight performance improvements over the LLM baseline in most cases, which provides potential for using LLMs in RDB-to-graph transformation.
@@
[…]
Table 1: Performance of ten RDB-to-graph modeling methods on the rel-f1 dataset under varying budget levels. Refer to Appendix D.1 for results on other datasets.
Task Name driver-dnf (AUC-ROC (%) $\uparrow$) driver-top3 (AUC-ROC (%) $\uparrow$) driver-position (MAE $\downarrow$)
Methods Budget (%) Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 74.557 74.557 74.557 74.557 74.557 81.879 81.879 81.879 81.879 81.879 3.8311 3.8311 3.8311 3.8311 3.8311
Random 73.225 73.450 73.592 73.745 73.755 79.627 80.165 80.425 80.438 80.604 3.8498 3.8435 3.8420 3.8405 3.8399
AR2N 73.140 73.140 73.140 73.140 73.140 78.106 78.106 78.106 78.106 78.106 3.9125 3.9125 3.9125 3.9125 3.9125
GF 73.461 74.557 74.557 74.557 74.557 81.879 81.879 81.879 81.879 81.879 3.8591 3.8409 3.8409 3.8409 3.8409
GB 73.254 74.040 74.394 74.394 74.394 80.148 80.558 80.558 80.558 80.558 3.8437 3.8437 3.8437 3.8437 3.8437
GL 73.023 73.645 73.726 73.775 73.775 79.172 79.292 79.299 79.299 79.299 3.8649 3.8512 3.8500 3.8500 3.8500
EA 73.082 73.393 73.590 73.786 73.929 79.153 79.774 80.116 80.273 80.549 3.8470 3.8451 3.8423 3.8394 3.8394
BO 73.328 73.624 73.837 73.919 74.070 79.424 79.799 80.127 80.152 80.348 3.8520 3.8404 3.8399 3.8394 3.8388
RL 73.332 73.526 73.602 73.657 73.986 79.043 79.385 79.439 79.906 80.084 3.8542 3.8449 3.8417 3.8411 3.8411
LLM 73.857 73.857 73.857 73.857 73.857 80.337 80.536 80.536 80.607 80.607 3.8534 3.8534 3.8500 3.8500 3.8500
LLM-CoT 73.338 73.574 73.882 73.882 73.882 78.660 78.693 78.693 78.817 78.817 3.8437 3.8437 3.8437 3.8437 3.8437
@@ 6 Conclusions
[…]
Acknowledgements
This work was partly supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00406985, 20%). This work was partly supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00438638, EntireDB2AI: Foundations and Software for Comprehensive Deep Representation Learning and Prediction on Entire Relational Databases, 50%) (No. RS-2024-00457882, AI Research Hub Project, 10%) (RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale, 10%) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST), 10%).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We reflect all the contributions and scope in the abstract and introduction(Section  1). In detail, we provide 3 contributions while describing the main claim of our paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We describe our limitations in Appendix E.1.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: Our paper doesn’t include any theoretical results.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: Our results are reproducible because all experiment results are computed from RDB2G-Bench, which we provide in this paper. We also release our datasets and codes in public (Github, Huggingface).
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We provide all the codes and datasets through RDB2G-Bench. We also construct documents to faithfully reproduce the main experimental results. You can check all the code and data from the following URLs:
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We mentioned our hyperparameter settings in Section 3.2 and Section 5.1. Also all the details are described in Appendix A.3 and Appendix C.1.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We report standard deviations in Figure 1 and error bars in Figure 2, and leverage box plots in Figure 3 to justify the statistical significance of our experiments.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We describe all the computer resources and time of execution in Section 3.2 (See Machines.).
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We strictly follow the NeurIPS Code of Ethics, particularly the source of our dataset, which RelBench has already anonymized all the data. Also, we don’t contain any privacy information in our datasets and codes.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We discuss various broader impacts of our paper in Appendix E.2.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We believe we do not have risks of misusing.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: We used the datasets and models of RelBench [29] implementation, which mentioned several times in our paper. Also, there is no license violation in our paper.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We release RDB2G-Bench which includes datasets and codes. We provide the documentation to help users that can be found in our GitHub link: https://github.com/chlehdwon/RDB2G-Bench, and Hugging Face link: https://huggingface.co/datasets/kaistdata/RDB2G-Bench.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: We don’t use any crowdsourcing or human-related tasks in our paper.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: We do not utilize any crowdsourcing or human-related tasks.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [Yes]
  64. Justification: We implemented LLM-based method as a benchmark baseline, which motivated by AutoG [7] paper. The differences between AutoG are described in Section 5.1, and all the prompts that we used are in Appendix F.
@@ A.1 RelBench Detail
Avito is a leading online advertisement platform, providing a marketplace for users to buy and sell a wide variety of products and services, including real estate, vehicles, jobs, and goods. The Avito Context Ad Clicks dataset on Kaggle is part of a competition aimed at predicting whether an ad will be clicked based on contextual information. This dataset includes user searches, ad attributes, and other related data to help build predictive GNNs.
@@ A.2 GNN Implementation
In our implementation, we employ Equation (1) and Equation (2) for the $\mathrm{LocalMP}^{(k)}$ of Row2Node and Row2Edge modeling, respectively, to naturally extend our settings to transformer-based architectures. To fit the GPS within a single GPU, we reduce the number of sampled neighbors from 128 to 32.
@@ A.3 Hyperparameter Details
Table 3 shows our learning rate choices, which are only tuned for each task. For the other hyperparameter, all tasks used two-layer GNNs with a batch size of 512, 128 dimensions, and the Adam [22] optimizer.
Specifically, neighbor sampling counts for subgraph extraction are set to 128 when all tables are represented as nodes, and 12 when some tables are represented as edges. A lower sampling count is employed when tables are modeled as edges, since the two-layer GNN effectively captures paths longer than two hops, resulting in a significantly larger graph scale that cannot be processed by a single GPU.
@@ Appendix B Additional Results regarding Our Observations on the RDB2G-Bench Datasets
As shown in Figure 25, while cross-GNN correlations are relatively low for the two tasks on the rel-f1 dataset (driver-top3 and driver-position) due to high variances in the performances, for the other tasks, cross-GNN correlations are high. The results for user-repeat, user-visits, and user-ad-visit are presented in Figure 6 in the main paper.
Figure 25: Additional results regarding Obs 5. Spearman correlations between different predictive GNNs—GraphSAGE with sum and mean aggregation (denoted as sum and mean), GIN, and GPS—are generally high (above 0.7) across the remaining tasks.
@@ C.1 Implementation Details of Baselines
Bayesian Optimization. The Bayesian Optimization (BO) baseline employs an iterative search leveraging a surrogate Multi-Layer Perceptron (MLP) based on the BANANAS [37] implementation. The input embeddings represent graph structures converted into fixed-sized arrays based on selected edges.
@@ D.1 Performance Results on the RDB2G-Bench Datasets
In this section, we provide additional performance details of ten RDB-to-graph modeling methods, supplementing the summary presented in Table 1 and Figure 7(a) in the main paper. Details are presented in Tables 4 - 6 and Figure 26.
Table 4: Performance of ten RDB-to-graph modeling methods on the rel-event dataset under varying budget levels.
Task Name user-ignore (AUC-ROC (%) $\uparrow$) user-repeat (AUC-ROC (%) $\uparrow$) user-attendance (MSE $\downarrow$)
Methods Budget (%) Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 82.823 82.823 82.823 82.823 82.823 82.291 82.291 82.291 82.291 82.291 0.2375 0.2375 0.2375 0.2375 0.2375
Random 79.543 80.797 81.427 81.548 81.572 78.034 79.588 81.280 81.430 81.430 0.2446 0.2423 0.2396 0.2391 0.2389
AR2N 82.222 82.222 82.222 82.222 82.222 77.651 77.651 77.651 77.651 77.651 0.2445 0.2445 0.2445 0.2445 0.2445
GF 77.655 78.213 78.233 78.945 80.931 77.789 79.163 80.348 80.508 80.981 0.2447 0.2423 0.2399 0.2399 0.2385
GB 82.222 82.222 82.222 82.222 82.222 78.748 80.469 80.744 81.018 81.258 0.2432 0.2414 0.2409 0.2407 0.2400
GL 79.256 79.809 80.534 80.809 81.120 79.012 79.158 79.558 80.481 80.739 0.2423 0.2418 0.2415 0.2398 0.2391
EA 79.939 80.523 81.011 81.299 81.482 78.989 80.739 81.249 81.313 81.333 0.2425 0.2404 0.2395 0.2390 0.2389
BO 78.787 80.550 81.015 81.162 81.288 77.516 80.360 80.831 81.120 81.424 0.2449 0.2403 0.2396 0.2394 0.2392
RL 79.474 80.376 80.527 80.719 80.963 79.341 80.703 80.930 81.040 81.094 0.2474 0.2421 0.2412 0.2410 0.2409
LLM 80.541 80.541 81.146 81.751 81.751 78.841 79.537 80.857 80.941 80.941 0.2400 0.2397 0.2397 0.2397 0.2397
LLM-CoT 82.222 82.222 82.222 82.222 82.222 81.259 81.259 81.259 81.259 81.259 0.2394 0.2394 0.2387 0.2387 0.2387
Table 5: Performance of ten RDB-to-graph modeling methods on the rel-avito dataset under varying budget levels.
Task Name user-clicks (AUC-ROC (%) $\uparrow$) user-visits (AUC-ROC (%) $\uparrow$)
Methods Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 67.931 67.931 67.931 67.931 67.931 66.332 66.332 66.332 66.332 66.332
Random 66.594 66.856 67.098 67.151 67.302 64.444 65.169 65.521 65.708 65.877
AR2N 64.660 64.660 64.660 64.660 64.660 65.971 65.971 65.971 65.971 65.971
GF 66.978 67.096 67.096 67.096 67.096 66.318 66.318 66.318 66.318 66.318
GB 66.561 66.791 66.791 66.791 66.791 66.332 66.332 66.332 66.332 66.332
GL 65.770 66.183 66.468 66.468 66.483 64.700 65.090 65.118 65.124 65.124
EA 66.778 66.944 66.958 67.204 67.232 65.352 65.397 66.047 66.116 66.122
BO 66.388 67.069 67.170 67.442 67.584 64.823 65.004 65.380 65.924 66.191
RL 66.341 66.771 66.801 66.902 66.983 63.840 64.228 64.379 64.963 65.249
LLM 66.066 66.308 66.648 67.092 67.323 66.246 66.246 66.246 66.246 66.246
LLM-CoT 66.597 66.597 66.597 66.597 66.597 66.246 66.246 66.246 66.246 66.246
Task Name ad-ctr (MSE $\downarrow$) user-ad-visit (MAP (%) $\uparrow$)
Methods Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 0.0389 0.0389 0.0389 0.0389 0.0389 3.6816 3.6816 3.6816 3.6816 3.6816
Random 0.0399 0.0397 0.0396 0.0396 0.0396 3.2824 3.5017 3.6084 3.6468 3.6487
AR2N 0.0397 0.0397 0.0397 0.0397 0.0397 3.6610 3.6610 3.6610 3.6610 3.6610
GF 0.0402 0.0394 0.0394 0.0394 0.0394 3.6453 3.6453 3.6453 3.6453 3.6453
GB 0.0392 0.0392 0.0392 0.0392 0.0392 3.6610 3.6610 3.6610 3.6610 3.6610
GL 0.0412 0.0406 0.0406 0.0406 0.0406 2.2905 2.8499 3.0970 3.0989 3.1171
EA 0.0404 0.0400 0.0398 0.0395 0.0395 2.6709 3.3097 3.4353 3.4391 3.4432
BO 0.0399 0.0396 0.0394 0.0392 0.0392 3.0748 3.2248 3.4774 3.6241 3.6759
RL 0.0411 0.0406 0.0405 0.0403 0.0403 2.5915 2.8723 3.2453 3.2475 3.3514
LLM 0.0397 0.0397 0.0397 0.0397 0.0397 3.6610 3.6610 3.6610 3.6610 3.6610
LLM-CoT 0.0394 0.0394 0.0394 0.0394 0.0394 3.6610 3.6634 3.6655 3.6655 3.6679
Table 6: Performance of ten RDB-to-graph modeling methods on the rel-stack dataset (left) and the rel-trial dataset (right) under varying budget levels.
Task Name post-post-related (MAP (%) $\uparrow$) study-outcome (AUC-ROC (%) $\uparrow$)
Methods Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 12.040 12.040 12.040 12.040 12.040 70.913 70.913 70.913 70.913 70.913
Random 11.751 11.803 11.835 11.867 11.900 70.178 70.338 70.372 70.375 70.407
AR2N 10.823 10.823 10.823 10.823 10.823 68.091 68.091 68.091 68.091 68.091
GF 11.903 11.903 11.903 11.903 11.903 69.572 69.572 69.572 69.572 69.572
GB 11.165 11.165 11.165 11.165 11.165 69.413 69.413 69.413 69.413 69.413
GL 10.555 10.555 10.555 10.555 10.555 69.490 69.490 69.490 69.490 69.490
EA 11.786 11.897 11.946 11.954 11.954 70.639 70.771 70.772 70.772 70.772
BO 11.860 11.888 11.888 11.888 11.888 70.376 70.376 70.376 70.376 70.376
RL 11.706 11.776 11.826 11.863 11.863 69.959 69.959 69.959 69.959 69.959
LLM 11.111 11.386 11.386 11.386 11.386 68.556 68.556 68.556 68.556 68.556
LLM-CoT 10.964 10.964 10.964 10.964 10.964 68.608 68.608 68.608 68.608 68.608
Figure 26: Performance details of ten RDB-to-graph modeling methods on each of the 12 predictive tasks. The performances are computed under varying budget levels, corresponding to the number of graph models evaluated.
@@ D.2 Analysis of Evaluation Time
In this section, we present a detailed analysis of evaluation time costs for each RDB-to-graph modeling method, which complements the results summarized in Figure 7(b) in the main paper. Details are provided in Table 7.
@@
Random 319016.09 0.59 540705.24
Greedy Forward Search (GF) 104246.09 0.39 267297.67
Greedy Backward Search (GB) 216032.31 0.51 423592.76
Greedy Local Search (GL) 284439.64 0.34 836587.18
Evolutionary Algorithm (EA) 696925.57 132.62 5255.06
Bayesian Optimization (BO) 595853.52 1244.17 478.92
Reinforcement Learning (RL) 753865.31 17.66 42687.73
LLM baseline (LLM) 50861.62 3588.22 14.17
LLM-CoT baseline (CoT) 28847.97 2924.82 9.86
D.3 Analysis of LLM Interpretability
Our analysis shows that interpretability can be enhanced when the LLM-CoT baseline provides explicit reasoning traces for each selected graph transformation action. During the action planning phase, we prompted the LLM to explain its rationale behind each operation, allowing us to capture structured insights about how it interprets temporal and structural dependencies.
As shown below, the LLM did not merely select from predefined actions but offered justifications referencing temporal order and causal consistency. These intuitive rationales demonstrate that the LLM’s reasoning process is consistent with the underlying graph structure, helping us understand why each transformation was made.
<reasoning>
1. Current State Analysis:
- The last action restored the FK-PK relationship between results and races, which improved performance slightly
- This suggests that temporal relationships between tables are important for DNF prediction
- The key tables for DNF prediction appear to be results, qualifying, races, and drivers
2. Examining Remaining Relationships:
- The FK relationship between standings and drivers was removed
- A drivers past performance and championship position could correlate with their risk-taking behavior or car reliability
3. Proposed Action:
- Restore the FK-PK relationship between standings and drivers
- Rationale:
* Driver standings contain cumulative performance metrics (points, position, wins)
* The temporal nature of standings data could help capture driver form and reliability trends
* Similar to how restoring race relationships improved performance, driver standings relationship might capture important patterns
4. Expected Impact:
- Should improve models ability to capture driver-specific patterns over time
- Aligns with the positive results seen from restoring other temporal relationships
Therefore, I recommend using add_fk_pk_edge to restore the relationship between standings and drivers.
</reasoning>